OpenStudy (anonymous):

Can someone help me with these two problems? Write the equation in logarithmic form. (1/3)^3=1/27 Write the equation in exponential form. log3 1/9= -2

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\huge b^x=y\iff \log_b(y)=x\] you should be able to switch back and forth quickly

OpenStudy (misty1212):

for example \[10^3=1000\iff \log_{10}(1000)=3\]and \[2^{-4}=\frac{1}{16}\iff \log_2(\frac{1}{16})=-4\]

OpenStudy (anonymous):

Okay... that makes sense. @misty1212

OpenStudy (misty1212):

so \[\left(\frac{1}{3}\right)^3=\frac{1}{27}=\iff?\]

OpenStudy (anonymous):

log1/3 (1/27)=3 Right? @misty1212

OpenStudy (misty1212):

yes exactly

OpenStudy (misty1212):

now for the other one you have to translate the other way \[?\iff \log_3(\frac{1}{9})=-2\]

OpenStudy (misty1212):

did you get it or are you stuck?

OpenStudy (anonymous):

It would be 3^-2=1/9 correct? @misty1212

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

Wait that can't be right because it's not a choice... oops

OpenStudy (misty1212):

which is true right? since \[3^{-2}=\frac{1}{3^2}=\frac{1}{9}\]

OpenStudy (misty1212):

yes it is right

OpenStudy (anonymous):

Ohhh I see..

OpenStudy (anonymous):

Thanks so much @misty1212 you were so helpful! :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]